OpenStudy (anonymous):

slope of the line (-11,6) (12,-3)

OpenStudy (anonymous):

\[\huge (\frac{ 1 }{ 2 },\frac{ 3 }{ 2 })\]

OpenStudy (anonymous):

slope =\[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

Slope will be 9/23

OpenStudy (anonymous):

yes!!

OpenStudy (anonymous):

sorry i found out the midpoints!

OpenStudy (anonymous):

\[\huge- \frac{ 9 }{ 23 }\]

OpenStudy (texaschic101):

slope(m) = (y2 - y1)/(x2 - x1) set 1 (-11,6) set 2 (12,-3) m = (-3 -6) / (12 - (-11) m = - 9 / 23 kapil is correct.....slope is - 9/23